OpenStudy (anonymous):

A rectangular plot of land will be bounded on one side by an existing stone wall and on the other three sides by a new fence. This rectangular plot of land is to be enclosed by the stone wall and new fence and then divided into two plots by another piece of fencing joining the midpoints of the section of the stone wall and the opposite side. There is 450 feet of fencing available for the project. What are the dimensions of the rectangular plot with the largest area?

OpenStudy (anonymous):

make sense so far? 

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k, then we have 450 = 3Y + X

OpenStudy (anonymous):

we need to maximize area, so that means we need an equation for area, and we need to derive it, set it equal to zero, and solve for the unknown.

OpenStudy (anonymous):

Area = X*Y if 450=3y+x then x=450-3y plug that into our area equation: Area = X*Y Area = (450-3y)*Y A = 450y-3y^2 now take the derivative of the equation 450y-3y^2

OpenStudy (anonymous):

can you do this? or should I show you? I gotta go soon

OpenStudy (anonymous):

I got it now! Thanks so much!!!

OpenStudy (anonymous):

ur welcome ^_^